TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-08 23:32:00
subject: 2\18 Pt-1 ISS On Orbit Status 18-02-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
2\18 ISS On-Orbit Status 18-02-2002
Part 1 of 2

ISS On-Orbit Status 2/18/03

All ISS systems continue to function nominally, except as noted 
previously or below. Day 87 in space for Expedition 6.

During a Russian ground site (RGS) pass on Daily Orbit 2, FE-1 Nikolai
Budarin completed another session with the Russian MedOps cardio 
experiment MO-1 (Study of the Bioelectric Activity of the Heart at 
Rest), with CDR Ken Bowersox assisting as CMO (crew medical officer). 
[During the test, transmittal of the Gamma-1M ECG (electrocardiogram) 
signal to the ground was required for about 5-6 minutes.]

Later, Budarin performed repair work on the Russian Potok-150MK (150 
micron) air filter unit of the SOGS air revitalization subsystem in 
the SM, installing a new circuit card delivered by Progress M-47/10P, 
after the troubleshooting on 12/27/02 for a failed fan. The IFM 
(in-flight maintenance) was supported by tag-up with ground 
specialists.

FE-2/SO Don Pettit yesterday successfully removed and replaced the 
LAB1P5 RPCM (remote power controller module), which controls the VTR1 
video recorder in the Lab RWS (robotics work station). [This R&R was 
deferred from last week (2/12), when the ground, upon trying to 
perform a preemptive refresh of the RPCM, received an unexpected 
health flag. While the event was thought to be fully understood, 
engineers wanted to assess it completely before proceeding yesterday.]

Bowersox completed another EVARM (EVA radiation monitoring) badge 
reading, moving the four pouches from the IV-CPDS (intravehicular 
charged particle directional spectrometer) to the EVARM reader and 
later returning the badges to their former location. He then 
downloaded the data from the reader to the HRF (human research 
facility) PC for later downlink.

The Foot (foot/ground reaction forces during space flight) experiment
activities scheduled for Ken Bowersox for today and tomorrow are being
rescheduled. [Foot team has determined that data from Bowersox's last
session was recorded on only half of the channels. Some data are 
missing, and Sox's marker button presses were also not recorded. Since 
the problem is not understood yet, the activities were deferred.]

The upload of the new Rev. 3 (R3) software to the SSMMUs (solid state 
mass memory units) of the C&C MDMs (command & control computers) began 
this morning and is currently underway, having started with the 
standby C&C (#1).  [C&C1 was then swapped with the backup C&C MDM 
(C&C2), followed by uploadings for the backup INT-1 (internal systems) 
and GNC2 (guidance, navigation and control) MDMs. Actual transition of 
the computers is scheduled for tomorrow and Thursday. For 
transitioning two of the PCS (portable computer system) laptops, 
Bowersox assisted by swapping their HHDs (hard disk drives) with the 
new software, after "go" from MCC-H due to the sequential nature of 
the R3 transition. MCC-H also uplinked procedures for contingency 
communications configurations, involving manual TDRS tracking, to 
restore comm with MCC-H in case that unexpected C&C or GNC MDM 
transitions occur at certain times during the three-day R3 transition 
period.]

Don Pettit completed the daily routine maintenance of the SOSh life 
support system, incl. ASU toilet subsystem, as well as the preparation 
of the IMS inventory database for automated file import/export.

Budarin continued the disinfection and washing of the Russian SVO-ZV 
water supply system equipment started yesterday, using a hand pump and 
special receptacles with a disinfectant solution delivered on 
Progress-247/10P and water.

Pettit took the daily ppCO2 (carbon dioxide partial pressure) readings 
in Lab and SM with the CDMK (CO2 monitoring kit), after breakfast and 
again before dinner.

The Vozdukh CO2 (carbon dioxide) removal system failed overnight, but 
it is now again up and running. Failure signature was like the 
preceding one in late January, and suspected cause is "something with 
the electrical currents in the system". ppCO2 (CO2 partial pressure) 
in the Russian segment (RS) was back down at 3.5 mmHg this morning. 
The U.S. CDRA (CO2 removal assembly) was powered on last night when 
ppCO2 in the Lab reached 6 mmHg. It is presently still on, with ppCO2 
this morning down at 4.2 mmHg and decreasing.

IMV (intermodular ventilation) troubleshooting is finished, after the 
crew cleaned out the Lab starboard forward fan assembly, finding not 
much lint this time, but "plastic wrapping".

The two-man EMU (extravehicular mobility unit) suit donning/doffing
demonstration is being scheduled for the crew this week, after the
procedures were successfully completed on the ground last Friday 
(2/14).  Photographic records of the ground run will be uplinked to 
the crew.  [Purpose of the activity is to demonstrate that two 
crewmembers can conduct a contingency EVA in U.S. EMU equipment from 
the Joint Airlock without the aid of a third crewmate (IV). The demo 
by Bowersox and Pettit will involve donning and doffing of suits and 
helmets, including hose connections, but neither prebreathing nor suit 
pressurization.]

Development of troubleshooting procedures for the MSG (microgravity 
science glovebox) continues on the ground. [A special "tiger" team at 
POC (Payload Operations Center, Huntsville) has looked into the 
feasibility of using the MSG's MWA (maintenance work area) to operate 
on-orbit payloads during the time the MSG is inoperable. It was 
determined that it is not possible, for multiple reasons, to operate 
the CSLM (Coarsening in Solid-Liquid Mixtures), InSpace (Investigating 
the Structure of Paramagnetic Aggregates from Colloidal Emulsions) and 
PFMI (P{ore Formation and Migration Investigation) payloads in the 
MWA. Science has to wait for the MSG to be restored to full service.]

During inspection of the RED (resistive exercise device), the crew 
found that Flexpack canister #1009 is exhibiting noises indicating a 
little metal-on-metal contact. The crew recommends using the can as a 
spare. Issue is under investigation by the ground.

When performing maintenance of the TVIS (treadmill with vibration 
isolation and stabilization) over the weekend, the crew had to tighten 
two loose screws on the port and starboard brackets holding the 
treadmill in the SM pit. They also discovered that one screw on each 
bracket is missing. After analysis, ground engineers found no issue 
with the situation, giving the go-ahead for the crew to use the TVIS 
without any restrictions.

Yesterday's dynamic test of the SM ODU manifold-2 yaw thruster (RZ, 
plane II), which failed during the recent Progress 10P docking, showed 
no problems with the systems involved. Calculated delta-velocity and 
burn time of the test were nominal, and sensors reported no pressure 
irregularities in the entire thruster/manifold system. The problem may 
have been with signal transmission in the associated electronics and 
is being investigated further.

A smoke alarm and a dP/dt (pressure differential) alarm went off in 
the RS yesterday but were found to be false, after other indicators 
reported neither smoke nor pressure drop.

The recent underspeed performance of the Progress M-47/10P thrusters 
during the ISS reboost is still under investigation. [1.1 m/sec of the 
1.8 m/sec deficit in delta-V has been identified to be due to a known 
operator's error, and steps are being taken with documentation and 
procedures to prevent re-occurrence. The remaining 0.7 m/sec variance 
still unaccounted for. Propulsion designers may soon propose two 
dynamic tests of the Progress thrusters, one for each four-jet 
manifold, each one "like a small reboost".  In view of other current 
issues such as software upgrading, large Beta angles, etc., the need 
for these tests is being re-evaluated. A proposal may be forthcoming 
next Friday, 2/21.]

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.